Two vessels contain mixtures of alcohol and water. In one there is three times as much as alcohol as water; in other, five times as much as water as alcohol. How much much be drawn off from each to fill a third vessel which holds 7 gallons, in order that its contents may be half alcohol and half water.
Answer by stanbon(75887)

Two vessels contain mixtures of alcohol and water.
In one there is three times as much as alcohol as water;
3/4 alcohol
-------
in other, five times as much as water as alcohol.
1/6 alcohol
How much much be drawn off from each to fill a third vessel which holds 7 gallons, in order that its contents may be half alcohol and half water.
-----------
Quantity: x + y = 7 gallons
Alcohol::::0.75x+0.1667y = 0.50*7
---------------------------------
Multiply 1st by 75
Multiply 2nd by 100
-----
75x + 75y = 75*7
75x + 16.67y = 50*7
---
Substract 2nd from 1st and solve for "y":
58.33y = 25*7
y = 3 gallons (amt to draw from the 2nd vessel)
---
x = 4 gallons (amt to draw from the 1st vessel)
================
Cheers,
Stan H.